• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Boeglin

    Log entry time 16:07:29 on June28,2007

    Entry number 206911

    keyword=Shift Summary

    Events:

    8:10 radiator in, take data on 12C
    8:35 beam off for path change
    Hall in Controlled access
    Changed to new kinematics
    10:45 Problem with right Q1: fluctuating
    11:15 reset Q1 R seems to be stable now
    11:15 left arm floor mark camera can be moved again
    13:00 Q2/Q3 left arm are fluctuating set to 0
    16:00 reset to proper values after cycling



    Run overview:
    time Run # Events Radiator
    8:10 1610 391K in